OpenStudy (anonymous):

Which is the best estimate of x based on rounding the constants and coefficients in the equation to the nearest integer? 6.2x + 1.2 = 8.9

OpenStudy (anonymous):

6.2x + 1.2 = 8.9 6x + 1 = 9 6x = 8 x = 8/6 x = 1.33 x is rounded to 1

OpenStudy (dobby1):

@oadurrani plz do not just give the answer this site is to help people learn not just to give answers away so because of this I am obligated to report you

OpenStudy (dobby1):

and that is not quiet how you would do that question you do not round up the numbers before you find x you round afterwards